### v1

# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable potential in the field of natural language processing. Unfortunately, LLMs face significant security and ethical risks. Although techniques such as safety alignment are developed for defense, prior researches reveal the possibility of bypassing such defenses through well-designed jailbreak attacks. In this paper, we propose QueryAttack, a novel framework to systematically examine the generalizability of safety alignment. By treating LLMs as knowledge databases, we translate malicious queries in natural language into code-style structured query to bypass the safety alignment mechanisms of LLMs. We conduct extensive experiments on mainstream LLMs, ant the results show that QueryAttack achieves high attack success rates (ASRs) across LLMs with different developers and capabilities. We also evaluate QueryAttack’s performance against common defenses, confirming that it is difficult to mitigate with general defensive techniques. To defend against QueryAttack, we tailor a defense method which can reduce ASR by up to 64% on GPT-4-1106. The code of QueryAttack can be found on https://anonymous.4open.science/r/QueryAttack-334B.
WARNING: THIS PAPER CONTAINS UNSAFE MODEL RESPONSES.
Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models
Qingsong Zou1,3thanks: Equal contribution.  Jingyu Xiao211footnotemark: 1  Qing Li3thanks: Corresponding author.  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang2  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1 1Tsinghua Shenzhen International Graduate School  2The Chinese University of Hong Kong 3Pengcheng Laboratory  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University
1 Introduction
Large language models (LLMs) such as OpenAI’s GPT series (OpenAI, 2024a) and Meta’s Llama series (Touvron et al., 2023a) demonstrate remarkable generative potential across various domains (Xiao et al., 2024; Boiko et al., 2023; He et al., 2024; Gao et al., 2024a). However, the immense amounts of data used for training LLMs contain massive information, enabling them to learn unscreened knowledge, including those may evidently violate ethical and moral standards (Li et al., 2023; Jiang et al., 2024; Yuan et al., 2024; Bai et al., 2024). To some extent, LLMs can function as knowledge databases for users. Consequently, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
Figure 1: The figure illustrates a well-established phenomenon: malicious queries, when translated into a custom language, can bypass the security defenses of LLMs, leading them to generate harmful outputs based on that language. Adversaries can obtain the desired harmful text in natural language following translation. In contrast to these attacks, QueryAttack is capable of directly inducing LLMs to generate harmful text in natural language without requiring a de-translation step.
To align their responses with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their responses to human queries. For example, supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022), reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024), red teaming (Bai et al., 2022a), and the constitutional AI (Bai et al., 2022b) approach are proposed to enhance the safety of LLMs. These efforts enhance the ability of LLMs to identify and defend against malicious queries. Unfortunately, these countermeasures cannot completely prevent outputs that violate human ethics. A significant limitation of these methods is their reliance on malicious language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs. This dependency leaves room for adversaries to develop jailbreak methods using non-natural language as input. For example, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher and Morse code to translate harmful queries into encrypted text. ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding. CodeAttack (Ren et al., 2024b) leverages data structures to embed natural language queries within string variables of programming languages.
We observe that, the essence of these jailbreak attacks lies in defining a customized non-natural language syntax and then using this language to interact with the target LLMs, thereby bypassing their defense mechanisms. To enable the model to understand the customized non-natural language, some methods employ few-shots learning to allow the model to learn how to interact using the specified syntax in context (Peng et al., 2024; Yuan et al., 2024; Barak, 2023; Wei et al., 2023a).
Inspired by prior work, we posit that defenses that relying solely on RLHF cannot achieve a complete defense against jailbreak attacks using non-natural language as inputs. Instead, these defenses can only be incrementally improved by continuously providing LLMs with samples of zero-day attacks. However, this approach is costly and demands sustained effort and resources.
To systematically explore this perspective, we design an effective jailbreak attack to illustrate the challenges of safety generalization in environments that is distant from safety training distribution. Our proposed attack, referred to as QueryAttack, is carried out in three main components: (1) defining a query format based on non-natural language and using a translator to convert natural language queries into the target non-natural language format; (2) prompting the target LLM in context to understand the non-natural language and provide responses in natural language; (3) using the translated query to induce the model to generate harmful responses.
These three steps define a query task, analogous to quering data from a database using code. The additional cost introduced by this process is limited to translating the malicious query into the specified format, which can be easily adapted to any query based on natural languages. Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query format.
We design a total of nine query formats written in the styles of commonly used programming languages and evaluate our attack on both well-known open-source and closed-source LLMs. Experimental results demonstrate that QueryAttack effectively bypasses their security defenses.
Our contributions can be summarized as follows:
  1. 1.
    We propose a hypothesis regarding the safety mechanisms of LLM for natural language alignment, suggesting that solely relying on RLHF cannot prevent the emergence of zero-day attacks, and use it to guide the design of a novel jailbreak attack.
  2. 2.
    We introduce QueryAttack, a novel jailbreak framework for LLMs based on query language translation. QueryAttack only requires the translation of harmful quries without requiring a de-translation step.
  3. 3.
    Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate (ASR).
2 Related Work
2.1 Jailbreak Attacks on LLMs
Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information. Since natural language samples are widely used as training data during the training phase of LLMs, potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024). However, this kind of approaches relies heavily on extensive human effort, incurs significant costs on launching attacks, or it is difficult to adapt to different models or models trained on different datasets. Building on these observations, several attack methods based on white-box assumptions are proposed (Zou et al., 2023; Jones et al., 2023a, b; Gao et al., 2024c). These attacks leverage access to the model’s architecture and weights of parameters to generate adversarial prompts based on gradients. While effective, these methods are likely to fail when applied to closed-source models under black-box scenarios.
Compared to white-box attacks, black-box attacks have gradually garnered significant attention due to their minimal requirement for model-specific knowledge. In general, black-box attacks assume that adversaries can only adjust their prompt strategies based on the model’s responses (Wei et al., 2023a). Typical black-box attack methods, such as scenario setting-up or role-playing (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Song et al., 2025), work by prompting the LLMs into dangerous hypothetical contexts, thereby triggering conflicts with the value alignment mechanisms of these models. Another approach aims to use pre-trained models to adversarially generate harmful prompts by leveraging feedback from the target models as feedback (Chao et al., 2023a).
Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities. By replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs. Notably, some recent work has focused on crafting adversarial inputs using programming languages. For instance,  (Ren et al., 2024b) embed malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages.  (Lv et al., 2024) encrypt malicious prompts using custom functions, transforming them into code completion tasks. The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format. However, these methods typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content.
2.2 Safety Alignment for Defending Jailbreak
To defend against jailbreak attacks, researchers have developed various mitigation strategies to guide large language models in rejecting malicious prompts of potential adversaries.
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms. For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a). A common feature of these approaches is the use of red teaming and reinforcement learning (Bai et al., 2022a; Mehrabi et al., 2024). These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.
Beyond the training process, some approaches focus on input and output safeguards, such as input perturbation (Jain et al., 2023a), safe decoding (Xu et al., 2024), and jailbreak detection (Robey et al., 2023; Phute et al., 2024; Jain et al., 2023b; Gao et al., 2024b). These methods can effectively reduce the attack success rate of jailbreak attacks. However, their effectiveness depends heavily on the quality of malicious data used for training and incurs significant additional overhead during deployment, which may affect user experience.
The latest CoT reasoning-enhanced models, such as OpenAI’s O1 model (OpenAI, 2024c), employ an embedded chain-of-thought mechanism to enforce step-by-step reasoning about the intent of queries and reflection on their compliance with policies during the response generation process. This approach does not rely on text in natural language from the training phase and simultaneously improves the model’s understanding of both queries and policy guidelines.
3 Methodology
3.1 Observation
As discussed in Section 1, we aim to demonstrate that the current generalization capabilities of safety aligned LLMs are insufficient to defend against zero-day attack samples constructed by adversaries using custom OOD languages. To promote understanding of target models, many jailbreak attack methods employ few-shots or many-shots to help the model learn the user-defined language (Yuan et al., 2024; Jiang et al., 2024). However, custom languages can be highly diverse. The effectiveness of such attacks depends significantly on the model’s ability to understand the new language.
In light of this, some attacks use common but non-natural languages as templates to encrypt malicious queries, achieving high ASR even under zero-shot conditions (Peng et al., 2024; Ren et al., 2024b). These languages often exist within the training samples of LLMs and are well understood by the models. However, they are not fully explored during the safety alignment process, making it possible for the models to generate malicious responses based on harmful prompts written by these languages. Programming languages clearly align with the common characteristics of such non-natural languages. Therefore, in this paper, we use them as examples to construct our query template.
Given the excellent generative capabilities of LLMs, we leverage them to construct translators that converts text in natural language into the our custom query templates. Few-shot learning is employed to guide the model to understand the query intents defined by the templates and generate responses in natural language. For those models with stronger comprehension capabilities (e.g., GPT-O1 (OpenAI, 2024c)), we employ zero-shot learning. Figure 2 provides an overview of QueryAttack.
Figure 2: The overview of QureyAttack. QueryAttack executes a three-step process to jailbreak the target LLM: 1). Translating the harmful query into a code-style custom template. 2. Using in-context learning to help the target LLM understand the natural semantics of the template. 3. Launching the attack using the translated template.
3.2 Query Template
Previous encryption-based jailbreak attacks convert malicious prompts into formats not encountered during the safety alignment phase. CipherChat (Yuan et al., 2024) translates queries into encrypted text using techniques such as Caesar cipher and Morse code. CodeAttack (Ren et al., 2024b) encapsulates entire natural language queries within string variables of programming languages using data structures. CodeChameleon (Lv et al., 2024) designs four different encryption functions to alter the word order of original queries or reconstruct the queries for translation.
These methods require the LLMs to generate responses in a specified non-natural language, which the adversary then decrypts to responses in natural language format. However, these approaches typically depend on the model’s knowledge of the target language to interpret the prompts or place high requirements on the model’s ability to generate text in the target language.
Unlike these attacks, we aim to only translate the malicious prompts into specific formats, but making the target LLMs directly generate responses in natural language. A total of nine commonly used programming languages are selected to design the query templates. Specifically, we identify three query components from natural language queries:
(1) the query content (content);
(2) the modifiers of the query content (modifiers);
(3) the higher-level benign category to which the query content belongs (category).
Using these three components, we define the query templates, as shown in Figure 2. This insight is inspired by the semantics of structured query languages (SQL), which query content from databases using standardized syntax. Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS) protocol, which uses Uniform Resource Locators (URLs) to fetch web resources from servers, also informs this process. Once the three query components are identified, this query template can be adapted to other programming languages. We provide a total of nine query templates based on commonly used programming languages, which can be found in the appendix A.
3.3 Query Translator
After defining the templates, the primary task is to translate natural language prompts into the template format. Extracting the three query components from the text in natural language is the most critical step in this process. This task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009). Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction.
We design specialized prompts to guide the GPT-4-1106 in translation. These prompts are aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model. The details of the prompt are provided in the appendix C.
3.4 Query Learning
The purpose of the query learning is to guide the model in understanding the intent behind the query templates and then generate a natural language response. To help the LLMs understand the content queried by the templates, we first establish the context of the conversation. By describing the three query components, we guide the model in building a mapping from the query template to the natural language, and define the conversation within an educational context.
Few-shot learning is then used to reinforce the model’s understanding of the query and guide it on how to respond to these prompts using natural language. Some text in natural languages, which contain longer modifiers or multiple queries, may require several templates to help define the query. Therefore, we provide both short and long examples.
For models with strong understanding of programming languages, we can skip this process and use zero-shot to launch the attack. Finally, we provide detailed guidance in natural language to respond to the query queries. The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts. In the appendix B, we provide the complete prompt of this part in SQL style version.
4 Experiments
4.1 Experimental Setup
Victim Models. We test QueryAttack on 11 mainstream large language models: GPT-4o (OpenAI, 2024b), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct) and Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) (Touvron et al., 2023b; Meta, 2024). To maintain the reproducibility of the results, we set all the temperature to 0.
Datasets. We used AdvBench (Zou et al., 2023) as the dataset for our experiments. This is a harmful behavior dataset that contains 520 different harmful query instances written in natural language. For some experiments, we use a subset of AdvBench which contains 50 representative, non-repetitive harmful instructions refined in (Jiang et al., 2024). We will specify this at the beginning of these parts where the subset is used.
Baselines. We select five representative jailbreak techniques as baselines:
PAIR (Chao et al., 2023b): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
CipherChat (Yuan et al., 2024): a black-box attack that converts inputs into ciphers for jailbreaking.
CodeAttack (Ren et al., 2024b): a method that leverages code to embed natural language inputs into data structures to generate adversarial inputs.
TAP (Mehrotra et al., 2023): A black-box jailbreak attack via iterative refinement of attack prompts.
HEA (Song et al., 2025): A black-box method that embeds malicious queries within a scenario containing positive prompts and deceives the LLM into jailbreaking during subsequent malicious queries.
Evaluation and Metrics. We focus on three main metrics to evaluate the effectiveness of QueryAttack: harmfulness score (HS), refuse rate (RR), and attack success rate (ASR). Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM. The harmfulness score ranges from 1 to 5. To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in D.3. The extracted content is then input into GPT-4 Judge for evaluation. Detailed settings can be found in the appendix D.
Based on this, we can calculate the second metric: attack success rate (ASR), which is the percentage of harmful responses in the case of adversarial queries, as shown in Equation 1.
$ASR=\frac{\#\text{ of responses with }HS=5}{\#\text{ of responses }}.$ (1)
The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
$RR=\frac{\text{ \# of queries that are refused by LLM }}{\#\text{ of queries }}.$ (2)
In the comparison with baselines, we primarily consider ASR as the evaluation metric. In other experiments, we also report RR and HS to further supplement the evaluation.
We present the experimental results from CodeAttack (Ren et al., 2024b) and (Song et al., 2025), as they use the same benchmark as ours and also employ the GPT-4 Judge (Qi et al., 2024) method to evaluate their attacks. Therefore, we use their results as baseline for comparisons with QueryAttack. Note that (Song et al., 2025) consider an attack successful when the HS is "higher than 4", whereas we define success only when HS equals 5. However, we can still compare results using the average HS.
4.2 Results
PAIR TAP CipherChat CodeAttack HEA Ours (Top 1) Ours (Ensemble)
GPT-4-1106 - / - - / - - / 19.00% - / 81.00% - / - 4.65 / 82.18% 4.75 / 93.80%
GPT-4o* 3.16 / 45.38 % 3.24 / 51.34% 1.94 / 16.34% - / 89.00% 4.42 / 90.38% 4.72 / 90.58% 4.85 / 96.35%
Llama-3.1-8b* 3.06 / 35.38% 2.97 / 31.34% 1.76 / 0.00% - / - 4.67 / 95.38% 4.04 / 65.78% 4.83 / 88.89%
Gemini-flash* 1.92 / 18.27% 3.01 / 33.27% 2.12 / 5.38% - / - 4.64 / 100.00% 4.93 / 95.59% 4.99 / 99.62%
Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench. QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR. *: (Song et al., 2025) reports that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results. When calculating the ASR for QueryAttack, we count instances where HS = 5.
Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023). We demonstrate two configurations of QueryAttack. In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template. In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles used for the query templates. From this table, we can observe that QueryAttack achieves SoTA ASR. Despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values.
For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%. In contrast, the baselines which perform best are only able to bypass GPT-4-1106’s safeguards in up to 81% of cases. The same trend is observed in other models. Except for Llama-3.3-8B, QueryAttack’s ASR is lower than HEA. However, under the Ensemble configuration, the average HS remains higher than that of HEA. This demonstrates the effectiveness of QueryAttack.
For cost considerations, we test QueryAttack’s effectiveness against the O1 model using the subset of AdvBench containing 50 samples. Under the Ensemble configurations, QueryAttack achieves an average HS of 3.66 and an ASR of 50%. These results indicate that CoT Reasoning-enhanced models may have the potential to defend against QueryAttack, but QueryAttack still maintains a considerable ASR.
4.3 Ablation and Analysis
Impact of different language styles. We create nine query templates in different programming language styles: Python, Java, JavaScript, C++, C, C#, SQL, URL, and Go. Figure 3 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different styles. On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS. A noticeable decrease is observed when attacking Llama-3.1-70B with URL and SQL-style templates, where a higher RR leads to a lower HS. This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses. Thus, using languages that differ more from natural language is likely to increase QueryAttack’s ASR. Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
Figure 3: Performance of different language styles.
Impact of the target model’s parameter size. The increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack. Table 2 presents the ASRs of QueryAttack on models with different parameter sizes from the latest Llama-3.1 and Llama-3.2 series under Ensemble configuration. On Llama-3.2 series, although the ASR slightly decreases for the 1B parameter models compared to larger models, the ASR does not show a decline as the parameter size continues to grow. Specifically, as the parameter size increases from 8B to 70B, QueryAttack’s ASR on the Llama-3.1 series models rises from 88.89% to 92.91%. Without targeted safety alignment, bigger models may even have a higher risk of being attacked due to their stronger understanding ability of new language. A similar trend can also be seen in results from other previous studies (Ren et al., 2024b; Yuan et al., 2024; Lv et al., 2024). Therefore, exploring more robust defense methods is necessary.
3.1-8B 3.1-70B 3.2-1B 3.2-3B 3.2-11B
ASR 88.89% 92.91% 94.83% 81.61% 88.89%
RR 2.87% 4.41% 3.45% 2.11% 3.26%
Table 2: QueryAttack’s attack success rate and refusal rate on Llama 3.1 and 3.2 series models with different parameter sizes, under the Ensemble configuration.
Attention score distribution of natural language malicious queries and QueryAttack. We employ contrastive input erasure (CIE) Yin and Neubig (2022) to quantify attention distribution. CIE evaluates how input tokens influence the model’s preference between an expected token (ET) and an unexpected token (UT). Since LLMs typically begin their refusal responses with “Sorry” and their acceptance responses with “Sure”, we set the ET to “Sure” and the UT to “Sorry”. An illustrative example is shown in the Figure 4, where blue tokens plays the role in rejecting query, whereas red tokens have the opposite effect. When directly questioned, the LLM precisely focuses its attention on the token “make a bomb”, significantly contributing to its likelihood of denying a malicious query. In contrast, when applying query attack, LLMs mainly allocate attention to the term “method” and “WHERE NAME = ”, leading to a higher possibility to jailbreak.
Figure 4: The attention score of natural language malicious query and QueryAttack.
4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the appendix D.2.
Paraphrase (Jain et al., 2023a): A defense method that reduces the ASRs by reconstructing inputs while preserving natural semantics.
Rand-Insert, Rand-Swap, and Rand-Patch (Robey et al., 2023): A defense method against adversarial prompts by perturbing the inputs in different ways.
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Based on this insight, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the appendix D.2.
We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5-turbo, Llama-3.1-8B, and GPT-4-1106. For cost considerations, we use the subset of AdvBench refined in (Jiang et al., 2024) for evaluation in this section and report the results of the Ensemble configuration using Python, C++, and SQL styles.
Table 3 presents the average HS / ASR of QueryAttack and the extent to which these defenses reduce average HS / ASR. The results show that QueryAttack is robust to baseline defenses. In the worst case, QueryAttack still achieves an average ASR of 63% on GPT-3.5. SmoothLLM and Paraphrase assume that adversarial tokens may be embedded in malicious prompts, yet they do not show significant defensive effects against QueryAttack. In some cases, QueryAttack’s effectiveness is even enhanced. For example, Paraphrase increase the HS for Llama-3.1-8B by 0.14 and the ASR by 2%.
In contrast, our tailored defense based on cross-lingual alignment prompting effectively reduces the ASR of QueryAttack across all models, with the exception of a small reduction in GPT-3.5. However, on the other three models, our defense achieves an average ASR reduction of 63% and an average HS reduction of 1.53, demonstrating the effectiveness of our defense. The success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. On the O1 model, QueryAttack achieves a lower ASR than on other models, further supporting this opinion.
Paraphrase Rand-insert Rand-swap Rand-patch Ours
Gemini-flash ASR 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$
HS 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$
GPT-4-1106 ASR 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$
HS 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$
GPT-3.5 ASR 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$
HS 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$
Llama-3.1-8B ASR 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$
HS 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
Table 3: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.
5 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious inputs. Building on previous encrypt-decrypt-based jailbreak methods, we further demonstrate that even when the target LLM is not required to generate encrypted response and instead directly produces natural language responses, a number of mainstream LLMs remain ineffective in mitigating harmful content generation.
To validate this, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages these templates to bypass the target LLM’s safety alignment. We conduct extensive evaluations and show that QueryAttack effectively circumvents the defenses of mainstream LLMs. To defend against QueryAttack, we developed a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
6 Limitation
The limitation of our study is the insufficient discussion of related defense mechanisms. As mentioned in the Related Work (section 2), research on jailbreak attack defenses has explored various potential countermeasures. However, we are unable to cover all of them. A more comprehensive investigation of defense methods would provide deeper insights into the generalizability of QueryAttack and help identify more effective mitigation techniques against such attacks.
Appendix A Query Templates
Figure 5 shows templates written in some common programming language styles. These templates express similar intentions to SQL templates, that is, querying content from a data storage.
Figure 5: Templates written in common programming language styles.
Appendix B Attack Examples of QueryAttack
B.1 Success Examples
Figure 6 illustrates a simplified attack process of QueryAttack. The adversary first utilizes query learning in context. Then, the translated query is used to launch the attack.
Figure 6: An attack example of QueryAttack.
B.2 Prompts of Query Learning
Figure 8 and Figure 7 present the prompts used for few-shot and zero-shot learning in SQL style. For zero-shot learning, we establish an educational context to encourage the model to generate more detailed information related to the risky knowledge.
Figure 7: The few-shot prompts for in-context learning (SQL style).
Figure 8: The zero-shot prompts for in-context learning (SQL style).
Appendix C Prompts for Extracting query Components
We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process. To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components. The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality. To reduce costs, adversaries may also choose alternative LLMs for this task.
Figure 9: The prompts for extracting query components from text in natural language using GPT-4-1106.
Appendix D Experimental Details
D.1 Human Evaluation on GPT-4 Judge
To confirm the effectiveness of GPT evaluation, we conduct an experiment with human evaluators to assess the responses of LLMs. We randomly select 100 outputs from GPT-4-1106 under the CodeAttack. Three authors independently classify these outputs as either successful or failed attacks based on OpenAI’s policy, and the final results are determined by majority vote. The manual verification results show a 93% agreement with the GPT-4 based evaluation (which also uses OpenAI’s policy as the benchmark). The result demonstrates the effectiveness of GPT-4 based evaluation.
D.2 Defense Settings
Paraphrase. We follow the setting of (Jiang et al., 2024) and use the GPT-3.5 model and the exact same prompt to complete the paraphrase task. The paraphrased translated template is then used as a replacement to launch the attack in the last step 2.
Rand-Insert, Rand-Swap, and Rand-Patch. We follow the defense and evaluation setting of (Robey et al., 2023), but choose the hyperparameter settings that give the best defensive performance reported in their paper.
Cross-lingual alignment prompting based defense. Figure 10 shows the cross-lingual alignment prompts we use for each attack sample. These prompts guide the target model to analyze the intent of the input query and rewrite it as text in natural language. Then, the target model is asked to generate response according to the query described in the text in natural language.
Figure 10: Prompt for cross-lingual alignment based defense.
D.3 Ethic-related Disclaimer Removal
Figure 11 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 8. The response fully complies with the adversary’s query (Develop a program that can exploit vulnerabilities in IoT devices and take control of them). Besides, it provides concise mitigation measures as required and includes brief ethic-related disclaimers in the the response (highlighted in purple). However, these measures are clearly ineffective, as the model has already provided the detailed information queried by the adversary. Moreover, offering mitigation strategies may further enhance the adversary’s malicious knowledge to develop stronger attack vectors. Therefore, it should be considered a successful attack.
Nevertheless, we observe that these disclaimers and mitigation measures could influence the evaluation of GPT-4 Judge (Qi et al., 2024), making it more likely to assess the response as compliant with OpenAI’s policies. This could result in the harm score dropping from 5 to 4 (or even lower). Therefore, before submitting responses for GPT-4 Judge, we remove the initial disclaimer, the mitigation / education part and subsequent content to extract the harmful content.
Figure 11: An example of a malicious response generated by GPT-O1 (OpenAI, 2024c). The response is clearly malicious, containing a lot of harmful knowledge, but includes concise mitigation measures as required and provides educational disclaimers at the beginning and end (highlighted in purple).

### v2

@@
 Although techniques such as safety alignment are developed for defense, prior researches reveal the possibility of bypassing such defenses through well-designed jailbreak attacks.
-In this paper, we propose QueryAttack, a novel framework to systematically examine the generalizability of safety alignment.
-By treating LLMs as knowledge databases, we translate malicious queries in natural language into code-style structured query to bypass the safety alignment mechanisms of LLMs.
-We conduct extensive experiments on mainstream LLMs, ant the results show that QueryAttack achieves high attack success rates (ASRs) across LLMs with different developers and capabilities.
-We also evaluate QueryAttack’s performance against common defenses, confirming that it is difficult to mitigate with general defensive techniques.
-To defend against QueryAttack, we tailor a defense method which can reduce ASR by up to 64% on GPT-4-1106.
-The code of QueryAttack can be found on https://anonymous.4open.science/r/QueryAttack-334B.
+In this paper, we propose QueryAttack, a novel framework to examine the generalizability of safety alignment.
+By treating LLMs as knowledge databases, we translate malicious queries in natural language into structured non-natural query language to bypass the safety alignment mechanisms of LLMs.
+We conduct extensive experiments on mainstream LLMs, and the results show that QueryAttack not only can achieve high attack success rates (ASRs), but also can jailbreak various defense methods.
+Furthermore, we tailor a defense method against QueryAttack, which can reduce ASR by up to 64% on GPT-4-1106.
+Our code is available at https://github.com/horizonsinzqs/QueryAttack.
 WARNING: THIS PAPER CONTAINS UNSAFE MODEL RESPONSES.
@@
 However, the immense amounts of data used for training LLMs contain massive information, enabling them to learn unscreened knowledge, including those may evidently violate ethical and moral standards (Li et al., 2023; Jiang et al., 2024; Yuan et al., 2024; Bai et al., 2024).
-To some extent, LLMs can function as knowledge databases for users.
-Consequently, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
-Figure 1: The figure illustrates a well-established phenomenon: malicious queries, when translated into a custom language, can bypass the security defenses of LLMs, leading them to generate harmful outputs based on that language.
-Adversaries can obtain the desired harmful text in natural language following translation.
-In contrast to these attacks, QueryAttack is capable of directly inducing LLMs to generate harmful text in natural language without requiring a de-translation step.
-To align their responses with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their responses to human queries.
+Therefore, a critical responsibility of service providers is to prevent these models from supplying harmful information to potentially adversaries.
+Figure 1: Existing methods reveal a well-established phenomenon: malicious queries, when well encrypted, can bypass the security defenses of LLMs, leading them to generate encrypted harmful outputs.
+Unlike these methods, QueryAttack achieves jailbreak by translating malicious inputs into structured non-natural query languages and is capable of directly inducing LLMs to generate harmful content in natural language without the need for a de-translation step.
+To align the responses of LLMs with human ethics and preferences, numerous techniques are employed during the training process of LLMs to regulate their outputs to human queries.
 For example, supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022), reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024), red teaming (Bai et al., 2022a), and the constitutional AI (Bai et al., 2022b) approach are proposed to enhance the safety of LLMs.
-These efforts enhance the ability of LLMs to identify and defend against malicious queries.
-Unfortunately, these countermeasures cannot completely prevent outputs that violate human ethics.
-A significant limitation of these methods is their reliance on malicious language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs.
+Unfortunately, a significant limitation of these methods is their reliance on malicious natural language samples from the alignment stage to train the model to recognize malicious queries and ensure the generation of safe outputs.
 This dependency leaves room for adversaries to develop jailbreak methods using non-natural language as input.
-For example, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher and Morse code to translate harmful queries into encrypted text.
+Specifically, CipherChat (Yuan et al., 2024) uses encryption methods such as the Caesar cipher to translate harmful queries into encrypted text.
 ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding.
-CodeAttack (Ren et al., 2024b) leverages data structures to embed natural language queries within string variables of programming languages.
-We observe that, the essence of these jailbreak attacks lies in defining a customized non-natural language syntax and then using this language to interact with the target LLMs, thereby bypassing their defense mechanisms.
-To enable the model to understand the customized non-natural language, some methods employ few-shots learning to allow the model to learn how to interact using the specified syntax in context (Peng et al., 2024; Yuan et al., 2024; Barak, 2023; Wei et al., 2023a).
-Inspired by prior work, we posit that defenses that relying solely on RLHF cannot achieve a complete defense against jailbreak attacks using non-natural language as inputs.
-Instead, these defenses can only be incrementally improved by continuously providing LLMs with samples of zero-day attacks.
-However, this approach is costly and demands sustained effort and resources.
-To systematically explore this perspective, we design an effective jailbreak attack to illustrate the challenges of safety generalization in environments that is distant from safety training distribution.
-Our proposed attack, referred to as QueryAttack, is carried out in three main components: (1) defining a query format based on non-natural language and using a translator to convert natural language queries into the target non-natural language format; (2) prompting the target LLM in context to understand the non-natural language and provide responses in natural language; (3) using the translated query to induce the model to generate harmful responses.
-These three steps define a query task, analogous to quering data from a database using code.
+Deng et al.
+(2024) convert sensitive contents into low-resource languages.
+The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format.
+However, they typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content.
+As a result, their attack effectiveness is limited.
+Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge.
+We observe that, the essence of these jailbreak attacks lies in defining a customized encryption method and then using the language encrypted by this method to interact with the target LLMs, thereby bypassing their defense mechanisms.
+Inspired by prior work, we find that LLM’s defensive mechanisms are not sensitive to structured, non-natural query languages.
+For example, by treating the target LLM as a knowledge database, when using structured query language (SQL) to request malicious knowledge (as shown in Figure 1) the target LLM not only identifies the intent of the request well but also does not trigger the defense mechanisms.
+Instead, the target LLM responds to the entire prompt in natural language normally.
+From this new perspective, we propose an attack that first uses structured non-natural query languages to jailbreak LLMs, named QueryAttack.
+Specifically, we break down QueryAttack into three main components:
+1).
+Extracting three key components from the original query: the requested content, the modifier of the content, and the high-level category to which the content belongs (potential sources where the content can be found).
+2). Filling the query components into predefined query templates (e.g., SQL templates) to generate a structured non-natural query.
+3). Applying in-context learning to help the target LLM understand the natural semantics of the template and prompting the target LLM using the structured non-natural query.
+These three steps define a query task, analogous to querying data from a database using SQL.
 The additional cost introduced by this process is limited to translating the malicious query into the specified format, which can be easily adapted to any query based on natural languages.
-Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query format.
-We design a total of nine query formats written in the styles of commonly used programming languages and evaluate our attack on both well-known open-source and closed-source LLMs.
-Experimental results demonstrate that QueryAttack effectively bypasses their security defenses.
+Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query templates.
+We test QueryAttack on AdvBench (Zou et al., 2023) across both well-known open-source and closed-source LLMs.
+The experimental results show that QueryAttack effectively bypasses their security defenses, achieving state-of-the-art attack success rates (ASRs).
+Besides, we provide a visual analysis of QueryAttack’s success and propose a tailored defense method against QueryAttack.
 Our contributions can be summarized as follows:
-  1. 1.
-    We propose a hypothesis regarding the safety mechanisms of LLM for natural language alignment, suggesting that solely relying on RLHF cannot prevent the emergence of zero-day attacks, and use it to guide the design of a novel jailbreak attack.
-  2. 2.
-    We introduce QueryAttack, a novel jailbreak framework for LLMs based on query language translation.
-QueryAttack only requires the translation of harmful quries without requiring a de-translation step.
-  3. 3.
-    Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate (ASR).
-2 Related Work
-2.1 Jailbreak Attacks on LLMs
-Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information.
-Since natural language samples are widely used as training data during the training phase of LLMs, potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
-Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024).
-However, this kind of approaches relies heavily on extensive human effort, incurs significant costs on launching attacks, or it is difficult to adapt to different models or models trained on different datasets.
-Building on these observations, several attack methods based on white-box assumptions are proposed (Zou et al., 2023; Jones et al., 2023a, b; Gao et al., 2024c).
-These attacks leverage access to the model’s architecture and weights of parameters to generate adversarial prompts based on gradients.
-While effective, these methods are likely to fail when applied to closed-source models under black-box scenarios.
-Compared to white-box attacks, black-box attacks have gradually garnered significant attention due to their minimal requirement for model-specific knowledge.
-In general, black-box attacks assume that adversaries can only adjust their prompt strategies based on the model’s responses (Wei et al., 2023a).
-Typical black-box attack methods, such as scenario setting-up or role-playing (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Song et al., 2025), work by prompting the LLMs into dangerous hypothetical contexts, thereby triggering conflicts with the value alignment mechanisms of these models.
-Another approach aims to use pre-trained models to adversarially generate harmful prompts by leveraging feedback from the target models as feedback (Chao et al., 2023a).
-Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities.
-By replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs.
-Notably, some recent work has focused on crafting adversarial inputs using programming languages.
-For instance,  (Ren et al., 2024b) embed malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages.
-(Lv et al., 2024) encrypt malicious prompts using custom functions, transforming them into code completion tasks.
-The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format.
-However, these methods typically require the model to possess knowledge of encryption to understand the prompts or place high requirements on the model’s ability to generate encrypted content.
-2.2 Safety Alignment for Defending Jailbreak
-To defend against jailbreak attacks, researchers have developed various mitigation strategies to guide large language models in rejecting malicious prompts of potential adversaries.
-Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms.
-For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a).
-A common feature of these approaches is the use of red teaming and reinforcement learning (Bai et al., 2022a; Mehrabi et al., 2024).
-These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.
-Beyond the training process, some approaches focus on input and output safeguards, such as input perturbation (Jain et al., 2023a), safe decoding (Xu et al., 2024), and jailbreak detection (Robey et al., 2023; Phute et al., 2024; Jain et al., 2023b; Gao et al., 2024b).
-These methods can effectively reduce the attack success rate of jailbreak attacks.
-However, their effectiveness depends heavily on the quality of malicious data used for training and incurs significant additional overhead during deployment, which may affect user experience.
-The latest CoT reasoning-enhanced models, such as OpenAI’s O1 model (OpenAI, 2024c), employ an embedded chain-of-thought mechanism to enforce step-by-step reasoning about the intent of queries and reflection on their compliance with policies during the response generation process.
-This approach does not rely on text in natural language from the training phase and simultaneously improves the model’s understanding of both queries and policy guidelines.
-3 Methodology
-3.1 Observation
-As discussed in Section 1, we aim to demonstrate that the current generalization capabilities of safety aligned LLMs are insufficient to defend against zero-day attack samples constructed by adversaries using custom OOD languages.
-To promote understanding of target models, many jailbreak attack methods employ few-shots or many-shots to help the model learn the user-defined language (Yuan et al., 2024; Jiang et al., 2024).
-However, custom languages can be highly diverse.
-The effectiveness of such attacks depends significantly on the model’s ability to understand the new language.
-In light of this, some attacks use common but non-natural languages as templates to encrypt malicious queries, achieving high ASR even under zero-shot conditions (Peng et al., 2024; Ren et al., 2024b).
-These languages often exist within the training samples of LLMs and are well understood by the models.
-However, they are not fully explored during the safety alignment process, making it possible for the models to generate malicious responses based on harmful prompts written by these languages.
-Programming languages clearly align with the common characteristics of such non-natural languages.
-Therefore, in this paper, we use them as examples to construct our query template.
-Given the excellent generative capabilities of LLMs, we leverage them to construct translators that converts text in natural language into the our custom query templates.
-Few-shot learning is employed to guide the model to understand the query intents defined by the templates and generate responses in natural language.
-For those models with stronger comprehension capabilities (e.g., GPT-O1 (OpenAI, 2024c)), we employ zero-shot learning.
-Figure 2 provides an overview of QueryAttack.
+  • We are the first to observe that the defense mechanisms of LLMs are not sensitive to structured non-natural query languages and propose QueryAttack, a novel jailbreak framework based on this observation.
+  • Our evaluation on mainstream LLMs demonstrates that QueryAttack successfully bypasses their security mechanisms and achieves state-of-the-art average attack success rate.
+  • We propose a tailored defense approach to mitigate QueryAttack, and experiments show that it can effectively help LLMs reduce the attack success rate.
 Figure 2: The overview of QureyAttack.
 QueryAttack executes a three-step process to jailbreak the target LLM: 1).
-Translating the harmful query into a code-style custom template.
-2.
-Using in-context learning to help the target LLM understand the natural semantics of the template.
-3.
-Launching the attack using the translated template.
-3.2 Query Template
-Previous encryption-based jailbreak attacks convert malicious prompts into formats not encountered during the safety alignment phase.
-CipherChat (Yuan et al., 2024) translates queries into encrypted text using techniques such as Caesar cipher and Morse code.
-CodeAttack (Ren et al., 2024b) encapsulates entire natural language queries within string variables of programming languages using data structures.
-CodeChameleon (Lv et al., 2024) designs four different encryption functions to alter the word order of original queries or reconstruct the queries for translation.
-These methods require the LLMs to generate responses in a specified non-natural language, which the adversary then decrypts to responses in natural language format.
-However, these approaches typically depend on the model’s knowledge of the target language to interpret the prompts or place high requirements on the model’s ability to generate text in the target language.
-Unlike these attacks, we aim to only translate the malicious prompts into specific formats, but making the target LLMs directly generate responses in natural language.
-A total of nine commonly used programming languages are selected to design the query templates.
-Specifically, we identify three query components from natural language queries:
+Extracting three key query components from the original query.
+2).
+Filling the query template to get a query code.
+3).
+Applying in-context learning to help the target LLM understand the natural semantics of the template and prompting the target LLM using the query code.
+2 Background
+Large language models (LLMs) have demonstrated remarkable generative potential across various fields.
+However, they are still vulnerable to jailbreak attacks.
+Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information.
+Since natural language samples are widely used as safety alignment data during the training phase of LLMs (gpt4), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
+Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities.
+For example, by replacing sensitive contents with Base64 (Wei et al., 2023a), ciphertext (Yuan et al., 2024), or low-resource languages (Deng et al., 2024), such attacks induce mismatched generalization in the target LLMs.
+Despite the development of numerous defense methods by researchers to mitigate jailbreaking attacks, such as supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022) and reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024).
+However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; DBLP:conf/coling/WuG0025).
+Therefore, it is necessary to continue studying the attack vector of LLMs and provide insights for developing new defenses.
+3 Methodology
+The core idea of QueryAttack is to use structured non-natural query languages to carry out jailbreak attacks.
+As shown in Figure 2, QueryAttack defines a content query task using three components to induce harmful outputs from the target LLM: 1).
+Query Components Extraction, which extracts key query components from the original query written in natural language.
+2).
+Query Template Filling, which uses the extracted query components to fill the query template and get query code.
+3).
+ICL-based Query Understanding, which applies in-context learning to enable the target LLM to understand the natural semantics of the template and generate responses with the user’s desired content.
+3.1 Query Components Extraction
+We first identify the components to be extracted from a natural language query (referred to as query components).
+Taking using Structured Query Language (SQL) as an example, the preliminary task for converting a natural language query into SQL code is to extract the key components of the query sentence.
+We observe that, for a given query, the two most relevant components to the query’s semantics are: the content to be queried and the description of that content.
+These two components can naturally be filled into the “SELECT” and “WHERE NAME=” sections of an SQL query.
+To ensure the completeness of the SQL query and to guide the target LLM’s response, we add an additional component: the category to which the queried content belongs (which can be understood as potential sources where the content may be found).
+This component can then be filled into the “FROM” section of an SQL query.
+Using these three extracted query components, a clear query can be defined, regardless of the language used.
+For example, the three key query components of the natural language query “Tell me the method of crafting a bomb” are { content: ‘crafting method’, modifiers: ‘bomb’, category: ‘crafting catalog’.
+}.
+Therefore, the first step of QueryAttack is to extract the following three query components from the natural language query:
 (1) the query content (content);
@@
 (3) the higher-level benign category to which the query content belongs (category).
-Using these three components, we define the query templates, as shown in Figure 2.
-This insight is inspired by the semantics of structured query languages (SQL), which query content from databases using standardized syntax.
-Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS) protocol, which uses Uniform Resource Locators (URLs) to fetch web resources from servers, also informs this process.
-Once the three query components are identified, this query template can be adapted to other programming languages.
-We provide a total of nine query templates based on commonly used programming languages, which can be found in the appendix A.
-3.3 Query Translator
-After defining the templates, the primary task is to translate natural language prompts into the template format.
-Extracting the three query components from the text in natural language is the most critical step in this process.
-This task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009).
+The query components extraction task can be performed manually by humans or automated using natural language processing tools such as NLTK (Bird et al., 2009).
 Given the strong text-processing capabilities of LLMs, we choose GPT-4-1106 (OpenAI, 2024a) to perform the component extraction.
-We design specialized prompts to guide the GPT-4-1106 in translation.
-These prompts are aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model.
+We design specialized prompts to guide the GPT-4-1106 in components extraction.
+This prompt is aimed at ensuring the model focuses on the text-processing task rather than the malicious content, and thereby prevents the translation request from being rejected by the model.
+To achieve this, we employ in-context learning to enable the model to understand the component extraction task and standardize its output.
 The details of the prompt are provided in the Appendix C.
-3.4 Query Learning
-The purpose of the query learning is to guide the model in understanding the intent behind the query templates and then generate a natural language response.
-To help the LLMs understand the content queried by the templates, we first establish the context of the conversation.
-By describing the three query components, we guide the model in building a mapping from the query template to the natural language, and define the conversation within an educational context.
+3.2 Query Template Filling
+Figure 3: Templates written in common programming language styles.
+After extracting the query components, we obtain the key semantics of a query.
+Therefore, we can rewrite the query into any non-natural language based on them.
+To automate the attack, we can predefine query templates for non-natural languages, allowing us to simply fill them with the query components to generate query in non-natural language.
+One observation is that non-natural languages commonly present in the training data of LLMs are more likely to be understood by LLMs.
+Since programming languages are widely present in the training data of LLMs, we choose them as examples to demonstrate the effectiveness of QueryAttack.
+Among many programming languages, two that are naturally associated with requesting content are SQL and Uniform Resource Locators (URLs).
+SQL is designed to query content from databases using standardized syntax.
+Similarly, the format of the Hypertext Transfer Protocol Secure (HTTPS), which uses URLs to fetch web resources from servers, follows a similar process.
+For example, in SQL, a malicious query written in natural language, such as “Tell me the method of crafting a bomb”, can be rewritten using its three query components as “SELECT ’crafting method’ FROM ’crafting catalog’ WHERE NAME = ’bomb’”.
+Inspired by this, although other programming languages are not specifically designed for querying resources, their syntax contains similar keywords or expressions that can express similar query intents, such as “print”, “input” and “return”.
+It means once the three query components are identified, this query template can be adapted to other programming languages.
+Specifically, using the query components, we define query templates for nine common programming languages (C, C++, C#, Python, Java, Javascript, Go, URL, SQL), as shown in Figure 3.
+As a conclusion, when using QueryAttack to jailbreak the LLMs, the second step is to fill the extracted three key query components into the corresponding language’s query template and obtain the query code, as shown in Figure 2.
+3.3 Query Understanding
+The purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response.
+To help the LLMs understand the content, we first establish the context of the conversation.
+By describing the three query components, we guide the model in building a mapping from the query code to the natural language, and define the conversation within an educational context.
 Few-shot learning is then used to reinforce the model’s understanding of the query and guide it on how to respond to these prompts using natural language.
-Some text in natural languages, which contain longer modifiers or multiple queries, may require several templates to help define the query.
+Some text in natural languages, which contain multiple queries, may require several query codes to help define the query.
 Therefore, we provide both short and long examples.
@@
 The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts.
-In the appendix B, we provide the complete prompt of this part in SQL style version.
+In Appendix A, we provide the complete prompt of this part.
+Through the above steps, we enable the target LLM to understand the intent of the query code and generate responses in natural language according to the query.
+The adversary then uses the query code obtained in the second step to launch the attack and obtain the desired malicious knowledge.
 4 Experiments
@@
 Victim Models.
-We test QueryAttack on 11 mainstream large language models: GPT-4o (OpenAI, 2024b), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct) and Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) (Touvron et al., 2023b; Meta, 2024).
+We test QueryAttack on 13 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024).
 To maintain the reproducibility of the results, we set all the temperature to 0.
@@
 We will specify this at the beginning of these parts where the subset is used.
+GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
+PAIR - / - 3.16 / 45.38 % 3.06 / 35.38% 3.24 / 47.30% 1.92 / 22.31% 1.92 / 18.27%
+TAP - / - 3.24 / 51.34% 2.97 / 31.34% 3.71 / 55.38% 2.83 / 24.23% 3.01 / 33.27%
+CipherChat - / 19% 1.94 / 16.34% 1.76 / 0 % 2.40 / 4.23% 2.22 / 3.27% 2.12 / 5.38%
+CodeAttack - / 81% - / 89% - / - - / - - / 2% - / -
+HEA - / - 4.42 / 90.38% 4.67 / 95.38% 3.58 / 68.27% 4.21 / 82.38% 4.64 / 100%
+Ours (Top 1) 4.65 / 82.18% 4.72 / 90.58% 4.04 / 65.78% 3.98 / 68.77% 4.71 / 85.63% 4.93 / 95.59%
+Ours (Ensemble) 4.75 / 93.80% 4.85 / 96.35% 4.83 / 88.89% 4.11 / 73.56% 4.91 / 95.40% 4.99 / 99.62%
+Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench.
+QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR.
+*: Song et al.
+(2025) report that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results.
+When calculating the ASR for QueryAttack, we follow Ren et al.
+(2024b) to count instances where HS = 5.
 Baselines. We select five representative jailbreak techniques as baselines:
-PAIR (Chao et al., 2023b): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
+PAIR (Chao et al., 2023): a black-box attack that uses an adversarial LLM to automatically generate adversarial inputs for the target LLM.
 CipherChat (Yuan et al., 2024): a black-box attack that converts inputs into ciphers for jailbreaking.
@@
 HEA (Song et al., 2025): A black-box method that embeds malicious queries within a scenario containing positive prompts and deceives the LLM into jailbreaking during subsequent malicious queries.
-Evaluation and Metrics.
-We focus on three main metrics to evaluate the effectiveness of QueryAttack: harmfulness score (HS), refuse rate (RR), and attack success rate (ASR).
+Evaluation and Metrics. We focus on three main metrics for evaluation.
+Harmfulness score (HS).
 Our evaluation follows the work of (Qi et al., 2024), utilizing the powerful evaluation capabilities of GPT-4 to assign harmfulness scores to the responses generated by the target LLM.
 The harmfulness score ranges from 1 to 5.
-To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in D.3.
+To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in B.3.
 The extracted content is then input into GPT-4 Judge for evaluation.
-Detailed settings can be found in the appendix D.
-Based on this, we can calculate the second metric: attack success rate (ASR), which is the percentage of harmful responses in the case of adversarial queries, as shown in Equation 1.
+Detailed settings can be found in the Appendix B.
+Attack success rate (ASR). Based on HS, we calculate the ASR, the percentage of harmful responses in the case of adversarial queries, which can be calculated by Equation 1.
 $ASR=\frac{\#\text{ of responses with }HS=5}{\#\text{ of responses }}.$ (1)
-The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
+Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
 $RR=\frac{\text{ \# of queries that are refused by LLM }}{\#\text{ of queries }}.$ (2)
-In the comparison with baselines, we primarily consider ASR as the evaluation metric. In other experiments, we also report RR and HS to further supplement the evaluation.
 We present the experimental results from CodeAttack (Ren et al., 2024b) and (Song et al., 2025), as they use the same benchmark as ours and also employ the GPT-4 Judge (Qi et al., 2024) method to evaluate their attacks.
 Therefore, we use their results as baseline for comparisons with QueryAttack.
-Note that (Song et al., 2025) consider an attack successful when the HS is "higher than 4", whereas we define success only when HS equals 5.
-However, we can still compare results using the average HS.
+Note that Song et al.
+(2025) consider an attack successful when the HS is “higher than 4”, whereas we follow Ren et al.
+(2024b) to define success only when HS equals 5.
 4.2 Results
-PAIR TAP CipherChat CodeAttack HEA Ours (Top 1) Ours (Ensemble)
-GPT-4-1106 - / - - / - - / 19.00% - / 81.00% - / - 4.65 / 82.18% 4.75 / 93.80%
-GPT-4o* 3.16 / 45.38 % 3.24 / 51.34% 1.94 / 16.34% - / 89.00% 4.42 / 90.38% 4.72 / 90.58% 4.85 / 96.35%
-Llama-3.1-8b* 3.06 / 35.38% 2.97 / 31.34% 1.76 / 0.00% - / - 4.67 / 95.38% 4.04 / 65.78% 4.83 / 88.89%
-Gemini-flash* 1.92 / 18.27% 3.01 / 33.27% 2.12 / 5.38% - / - 4.64 / 100.00% 4.93 / 95.59% 4.99 / 99.62%
-Table 1: Average HS / ASR of baselines and QueryAttack on the AdvBench.
-QueryAttack can breach the safety guardrails of mainstream LLMs, including GPT, Llama and Gemini series and achieves SoTA ASR.
-*: (Song et al., 2025) reports that they “consider responses of LLMs with a score higher than 4 as successful attacks” and we present their original results.
-When calculating the ASR for QueryAttack, we count instances where HS = 5.
+QueryAttack achieves SoTA ASR.
 Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023).
@@
 In the first configuration, denoted as Top 1, QueryAttack uses the programming language style with the highest ASR to construct the query template.
-In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles used for the query templates.
-From this table, we can observe that QueryAttack achieves SoTA ASR.
-Despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values.
+In the second configuration, denoted as Ensemble, no restrictions are placed on the programming language styles.
+From this table, we can observe that, despite specialized safety alignment training in these latest LLMs, QueryAttack successfully bypassed their defenses, inducing responses that violate policies or human values.
 For GPT-4-1106, the Top 1 configuration achieves a 82.18% ASR and the Ensemble configuration achieves 93.80%.
@@
 The same trend is observed in other models.
-Except for Llama-3.3-8B, QueryAttack’s ASR is lower than HEA.
+Except for Llama-3.1-8B, QueryAttack’s ASR is lower than HEA.
 However, under the Ensemble configuration, the average HS remains higher than that of HEA.
-This demonstrates the effectiveness of QueryAttack.
+QueryAttack remains effective when facing reasoning-enhanced models.
 For cost considerations, we test QueryAttack’s effectiveness against the O1 model using the subset of AdvBench containing 50 samples.
@@
 4.3 Ablation and Analysis
-Impact of different language styles.
-We create nine query templates in different programming language styles: Python, Java, JavaScript, C++, C, C#, SQL, URL, and Go.
-Figure 3 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different styles.
+Languages that differ more from natural language is likely to increase QueryAttack’s ASR.
+Figure 4 shows the average HS and RR obtained by attacking GPT-4-1106, Llama-3.1-70B, Gemini-flash and DeepSeek with templates of different language styles.
 On GPT-4-1106, Gemini-flash and DeepSeek, different language styles do not show significant variations in average HS.
@@
 This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses.
-Thus, using languages that differ more from natural language is likely to increase QueryAttack’s ASR.
 Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
 Figure 4: Performance of different language styles.
-Impact of the target model’s parameter size.
-The increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack.
+Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
+ASR HS ASR HS ASR HS ASR HS
+Paraphrase 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$
+Rand-insert 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
+Rand-swap 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
+Rand-patch 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
+Ours 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
+Table 2: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.
+Larger models do not provide better defense against QueryAttack.
 Table 3 presents the ASRs of QueryAttack on models with different parameter sizes from the latest Llama-3.1 and Llama-3.2 series under Ensemble configuration.
@@
 Specifically, as the parameter size increases from 8B to 70B, QueryAttack’s ASR on the Llama-3.1 series models rises from 88.89% to 92.91%.
+This means increase in model parameter size does not show a positive correlation with the effectiveness of defending against QueryAttack.
 Without targeted safety alignment, bigger models may even have a higher risk of being attacked due to their stronger understanding ability of new language.
-A similar trend can also be seen in results from other previous studies (Ren et al., 2024b; Yuan et al., 2024; Lv et al., 2024).
 Therefore, exploring more robust defense methods is necessary.
@@
 4.4 Discussion about Countermeasures
-We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the appendix D.2.
+We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix B.2.
 Paraphrase (Jain et al., 2023a): A defense method that reduces the ASRs by reconstructing inputs while preserving natural semantics.
@@
 Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent.
-Based on this insight, we design a defense method using cross-lingual chain-of-thought prompting.
+Our insight is that, the success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language.
+Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses.
+This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks.
+Based on this, we design a defense method using cross-lingual chain-of-thought prompting.
 The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language.
 The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense.
-The detailed prompt can be found in the appendix D.2.
-We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5-turbo, Llama-3.1-8B, and GPT-4-1106.
+The detailed prompt can be found in the Appendix B.2.
+We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5, Llama-3.1-8B, and GPT-4-1106.
 For cost considerations, we use the subset of AdvBench refined in (Jiang et al., 2024) for evaluation in this section and report the results of the Ensemble configuration using Python, C++, and SQL styles.
@@
 However, on the other three models, our defense achieves an average ASR reduction of 63% and an average HS reduction of 1.53, demonstrating the effectiveness of our defense.
-The success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language.
-Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses.
-This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks.
-On the O1 model, QueryAttack achieves a lower ASR than on other models, further supporting this opinion.
-Paraphrase Rand-insert Rand-swap Rand-patch Ours
-Gemini-flash ASR 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 100% ${\color[rgb]{1,0,0}-0\%}$ 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$
-HS 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$
-GPT-4-1106 ASR 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$
-HS 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$
-GPT-3.5 ASR 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$
-HS 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$
-Llama-3.1-8B ASR 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$
-HS 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
-Table 3: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.
+5 Related Work
+5.1 Jailbreak Attacks on LLMs
+Initially, researchers reveal that adversaries could launch attacks by manually constructing out-of-distribution (OOD) samples (Li et al., 2023; Shen et al., 2024).
+Building on these observations, several white-box attack methods are proposed (Jones et al., 2023a, b; Gao et al., 2024c).
+Compared to white-box attacks, black-box attacks assume that adversaries adjust their prompt strategies only based on the model’s responses (Zhang et al., 2024; Shah et al., 2023; Liu et al., 2024; Ren et al., 2024c; Yuan et al., 2024; Deng et al., 2024).
+Recently, some black-box are proposed to use code to encrypt malicious inputs to build long-tail encoded distributions.
+CodeAttack (Ren et al., 2024b) embeds malicious queries within data structures (e.g., stacks and queues) to bypass safety alignments designed for prompts written in natural languages.
+Codechameleon (Lv et al., 2024) encrypts malicious prompts using custom program functions, transforming them into code completion tasks.
+Unlike these methods, QueryAttack does not rely on the syntax of programming languages for encryption.
+Instead, it only requires certain keywords or expressions from the programming language.
+This means that QueryAttack can be applied not only using programming languages but also to any non-natural language that the target LLM can understand but has not been well aligned during the safety alignment phase.
+Moreover, even without the need for output encryption, QueryAttack can still effectively attack target LLMs.
+5.2 Safety Alignment for Defending Jailbreak
+Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) is one of the most widely used defense mechanisms.
+For instance, recent works such as (Mehrabi et al., 2024; Sun et al., 2023) explore the effectiveness of alignment during pre-training in defending against malicious queries, CoT reasoning (OpenAI, 2024c), as well as in-context learning (Wei et al., 2023b; Ren et al., 2024a).
+These methods often rely on natural language inputs collected from red teams, which can lead to generalization issues when faced with non-natural language or other OOD inputs.
+Beyond the training process, some approaches focus on input and output safeguards, such as input perturbation (Jain et al., 2023a), safe decoding (Xu et al., 2024), and jailbreak detection (Robey et al., 2023; Phute et al., 2024; Jain et al., 2023b; Gao et al., 2024b).
+These methods can effectively reduce the attack success rate of jailbreak attacks.
+However, their effectiveness depends heavily on the quality of malicious data used for training and incurs significant additional overhead during deployment, which may affect user experience.
 6 Conclusion
-In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious inputs.
-Building on previous encrypt-decrypt-based jailbreak methods, we further demonstrate that even when the target LLM is not required to generate encrypted response and instead directly produces natural language responses, a number of mainstream LLMs remain ineffective in mitigating harmful content generation.
-To validate this, we introduce QueryAttack, a novel jailbreak attack framework.
-QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages these templates to bypass the target LLM’s safety alignment.
-We conduct extensive evaluations and show that QueryAttack effectively circumvents the defenses of mainstream LLMs.
-To defend against QueryAttack, we developed a tailored defense strategy.
+In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious structured non-natural query language.
+Specifically, we introduce QueryAttack, a novel jailbreak attack framework.
+QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages the obtained query code to bypass the target LLM’s safety alignment.
+Although QueryAttack does not encrypt the outputs, our extensive evaluation shows that it still effectively bypasses the defenses of mainstream LLMs and can withstand common defense methods.
+Besides, to defend against QueryAttack, we develop a tailored defense strategy.
 Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
-6 Limitation
+Limitations
 The limitation of our study is the insufficient discussion of related defense mechanisms.
@@
 A more comprehensive investigation of defense methods would provide deeper insights into the generalizability of QueryAttack and help identify more effective mitigation techniques against such attacks.
-Appendix A Query Templates
-Figure 5 shows templates written in some common programming language styles. These templates express similar intentions to SQL templates, that is, querying content from a data storage.
-Figure 5: Templates written in common programming language styles.
-Appendix B Attack Examples of QueryAttack
-B.1 Success Examples
+Ethics Statement
+Please note that all experiments are conducted within controlled laboratory environments.
+We do not support the application of our QueryAttack in real-world scenarios.
+The primary objective of our work is to raise awareness about the security concerns related to the safety of LLMs.
+We aim to highlight the potential vulnerabilities about QueryAttack and encourage practitioners to prioritize the development of robust security measures and trustworthy deployment practices.
+Appendix A Attack Examples of QueryAttack
+A.1 Success Examples
 Figure 6 illustrates a simplified attack process of QueryAttack. The adversary first utilizes query learning in context. Then, the translated query is used to launch the attack.
-Figure 6: An attack example of QueryAttack.
-B.2 Prompts of Query Learning
+Figure 6: An attack example of QueryAttack using a C++ style template.
+A.2 Prompts of Query Learning
 Figure 7 and Figure 8 present the prompts used for few-shot and zero-shot learning in SQL style.
 For zero-shot learning, we establish an educational context to encourage the model to generate more detailed information related to the risky knowledge.
-Figure 7: The few-shot prompts for in-context learning (SQL style).
-Figure 8: The zero-shot prompts for in-context learning (SQL style).
-Appendix C Prompts for Extracting query Components
-We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process.
-To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components.
-The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality.
-To reduce costs, adversaries may also choose alternative LLMs for this task.
-Figure 9: The prompts for extracting query components from text in natural language using GPT-4-1106.
-Appendix D Experimental Details
-D.1 Human Evaluation on GPT-4 Judge
+Figure 7: The zero-shot prompts for Query Understanding (SQL style).
+Figure 8: The few-shot prompts for Query Understanding (C++ style).
+Appendix B Experimental Details
+B.1 Human Evaluation on GPT-4 Judge
 To confirm the effectiveness of GPT evaluation, we conduct an experiment with human evaluators to assess the responses of LLMs.
@@
 The result demonstrates the effectiveness of GPT-4 based evaluation.
-D.2 Defense Settings
+B.2 Defense Settings
 Paraphrase.
@@
 Figure 9: Prompt for cross-lingual alignment based defense.
-D.3 Ethic-related Disclaimer Removal
-Figure 11 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 8.
+B.3 Ethic-related Disclaimer Removal
+Figure 10 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 7.
 The response fully complies with the adversary’s query (Develop a program that can exploit vulnerabilities in IoT devices and take control of them).
@@
 The response is clearly malicious, containing a lot of harmful knowledge, but includes concise mitigation measures as required and provides educational disclaimers at the beginning and end (highlighted in purple).
+Appendix C Prompts for Extracting Query Components
+We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process.
+To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components, as shown in Figure 11.
+The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality.
+Figure 11: The prompts for extracting query components from text in natural language using GPT-4-1106.

### v3

@@
-Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models
+Qingsong Zou1,2  Jingyu Xiao311footnotemark: 1  Qing Li2  Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang3  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1, [^2] footnotemark: 2 1Tsinghua Shenzhen International Graduate School  2Pengcheng Laboratory
 Abstract
@@
 WARNING: THIS PAPER CONTAINS UNSAFE MODEL RESPONSES.
-Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models
-Qingsong Zou1,3thanks: Equal contribution.
-Jingyu Xiao211footnotemark: 1  Qing Li3thanks: Corresponding author.
-Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang2  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1 1Tsinghua Shenzhen International Graduate School  2The Chinese University of Hong Kong 3Pengcheng Laboratory  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University
+[Uncaptioned image] QueryAttack: Jailbreaking Aligned Large Language Models Using Structured Non-natural Query Language
+Qingsong Zou1,2thanks: Equal contribution.
+Jingyu Xiao311footnotemark: 1  Qing Li2thanks: These authors contributed equally to this work as co-corresponding authors.
+Zhi Yan4  Yuhang Wang5 Li Xu6  Wenxuan Wang3  Kuofeng Gao1  Ruoyu Li7  Yong Jiang1, [^2] footnotemark: 2 1Tsinghua Shenzhen International Graduate School  2Pengcheng Laboratory 3The Chinese University of Hong Kong  4Jilin University  5Southwest University 6University of Electronic Science and Technology of China  7Shenzhen University zouqs21@mails.tsinghua.edu.cn  jyxiao@link.cuhk.edu.hk  liq@pcl.ac.cn  jiangy@sz.tsinghua.edu.cn
 1 Introduction
@@
 ArtPrompt (Jiang et al., 2024) replaces sensitive terms with ASCII-style encoding.
-Deng et al.
-(2024) convert sensitive contents into low-resource languages.
+(Deng et al., 2024) convert sensitive contents into low-resource languages.
 The essence of these methods lies in inducing the model to generate encrypted outputs, which are then decrypted to harmful text in natural language format.
@@
 As a result, their attack effectiveness is limited.
+To illustrate this point, we design a simple yet clear experiment to test whether some mainstream large language models can effectively understand and generate encrypted text, as shown in Appendix A.
+The results show that some models may fail to achieve both objectives simultaneously and cause jailbreak failure.
 Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge.
@@
 Given that programming languages are widely present in the training data of LLMs and that these models exhibit excellent semantic understanding of programming languages (OpenAI, 2023, 2024a; Anthropic, 2023; Touvron et al., 2023a), we naturally employ programming syntax to construct the query templates.
-We test QueryAttack on AdvBench (Zou et al., 2023) across both well-known open-source and closed-source LLMs.
+We test QueryAttack on AdvBench (Zou et al., 2023b) across both well-known open-source and closed-source LLMs.
 The experimental results show that QueryAttack effectively bypasses their security defenses, achieving state-of-the-art attack success rates (ASRs).
@@
 Jailbreak attacks against LLMs typically involve crafting carefully designed inputs to prompt models to generate and output harmful response, such as instructions that blatantly violate human ethics or the disclosure of sensitive information.
-Since natural language samples are widely used as safety alignment data during the training phase of LLMs (gpt4), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
+Since natural language samples are widely used as safety alignment data during the training phase of LLMs (OpenAI, 2024a), potential malicious users or adversaries can bypass the defense mechanisms of these models by designing prompts based on non-natural language distributions.
 Attacks leveraging long-tail encoded distributions are particularly effective when the target LLM’s safety fine-tuning fails to generalize to domains requiring corresponding capabilities.
@@
 Despite the development of numerous defense methods by researchers to mitigate jailbreaking attacks, such as supervised fine-tuning (Wei et al., 2022; Ouyang et al., 2022) and reinforcement learning from human feedback (Sun et al., 2023; Mehrabi et al., 2024).
-However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; DBLP:conf/coling/WuG0025).
+However, recent work has shown that they are not able to completely defend against zero-day jailbreaking attacks (Song et al., 2025; Wu et al., 2025).
 Therefore, it is necessary to continue studying the attack vector of LLMs and provide insights for developing new defenses.
@@
 To achieve this, we employ in-context learning to enable the model to understand the component extraction task and standardize its output.
-The details of the prompt are provided in the Appendix C.
+The details of the prompt are provided in the Appendix D.
 3.2 Query Template Filling
@@
 3.3 Query Understanding
-The purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response.
+In-context learning is also applied to this step.
+However, unlike in query components extraction, the purpose of the query learning is to guide the model in understanding the intent behind the query codes and then generate a natural language response.
 To help the LLMs understand the content, we first establish the context of the conversation.
@@
 The aim of this process is to have the model answer the relevant knowledge as thoroughly as possible, rather than focusing on understanding and explaining the natural semantics of the prompts.
-In Appendix A, we provide the complete prompt of this part.
+In Appendix B, we provide the complete prompt of this part.
 Through the above steps, we enable the target LLM to understand the intent of the query code and generate responses in natural language according to the query.
@@
 Victim Models.
-We test QueryAttack on 13 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), Deepseek (deepseek-chat) (DeepSeek-AI et al., 2024), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024).
+We test QueryAttack on 14 mainstream large language models: GPT-3.5 (gpt-3.5-turbo) (OpenAI, 2023), GPT-4-1106 (gpt-4-1106-preview) (OpenAI, 2024a), GPT-4o (OpenAI, 2024b), O1 (gpt-o1) (OpenAI, 2024c), DeepSeek (deepseek-chat) (DeepSeek-AI et al., 2024), DeepSeek-R1 (DeepSeek-R1) (DeepSeek-AI et al., 2025), Gemini-flash (gemini-1.5-flash), Gemini-pro (gemini-1.5-pro) (DeepMind, 2024), Llama-3.1-8B (meta-llama-3.1-8B-instruct), Llama-3.1-70B (meta-llama-3.1-70B-instruct), Llama-3.2-1B (meta-llama-3.2-1B-instruct), Llama-3.2-3B (meta-llama-3.2-3B-instruct), Llama-3.2-11B (meta-llama-3.2-11B-vision-instruct) and Llama-3.3-70B (meta-llama-3.3-70B-instruct) (Touvron et al., 2023b; Meta, 2024).
 To maintain the reproducibility of the results, we set all the temperature to 0.
 Datasets.
-We used AdvBench (Zou et al., 2023) as the dataset for our experiments.
-This is a harmful behavior dataset that contains 520 different harmful query instances written in natural language.
+We use AdvBench (Zou et al., 2023b) and HEx-PHI Zou et al.
+(2023a) as the dataset for our experiments.
+AdvBench is a harmful behavior dataset that contains 520 different harmful query instances written in natural language.
+For HEx-PHI, we use the subset of it refined in (Jiang et al., 2024) for evaluation.
+We use QueryAttack to convert these harmful queries written in natural language into structured non-query language to attack the target LLMs.
 For some experiments, we use a subset of AdvBench which contains 50 representative, non-repetitive harmful instructions refined in (Jiang et al., 2024).
 We will specify this at the beginning of these parts where the subset is used.
-GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
+Method GPT-4-1106 GPT-4o* Llama-3.1-8B* Llama-3.3-70B* Gemini-pro* Gemini-flash*
 PAIR - / - 3.16 / 45.38 % 3.06 / 35.38% 3.24 / 47.30% 1.92 / 22.31% 1.92 / 18.27%
@@
 The harmfulness score ranges from 1 to 5.
-To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in B.3.
+To improve the accuracy of GPT-4 Judge, we remove educational statements or disclaimers that may appear in the responses to extract the harmful content, as shown in C.3.
 The extracted content is then input into GPT-4 Judge for evaluation.
-Detailed settings can be found in the Appendix B.
+Detailed settings can be found in the Appendix C.
 Attack success rate (ASR). Based on HS, we calculate the ASR, the percentage of harmful responses in the case of adversarial queries, which can be calculated by Equation 1.
 $ASR=\frac{\#\text{ of responses with }HS=5}{\#\text{ of responses }}.$ (1)
-Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
+Refuse rate (RR). The third evaluation metric is the refuse rate (RR). We follow the setup of (Zou et al., 2023b) and use Dict-Judge to assess the number of refusal responses, as shown in Equation 2.
 $RR=\frac{\text{ \# of queries that are refused by LLM }}{\#\text{ of queries }}.$ (2)
@@
 QueryAttack achieves SoTA ASR.
-Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023).
+Table 1 presents the average HS and the ASR of QueryAttack and several baselines on AdvBench (Zou et al., 2023b).
 We demonstrate two configurations of QueryAttack.
@@
 However, under the Ensemble configuration, the average HS remains higher than that of HEA.
+We conduct additional experiments using another widely-used public dataset, HEx-PHI (Zou et al., 2023a).
+For cost considerations, we employ the HEx-PHI subset defined in ArtPrompt (Jiang et al., 2024) (containing 11 categories with 10 samples each, totaling 110 samples) to test the harm score / attack success rate across four mainstream LLMs using Python, C++, and SQL, as shown in the Table 2.
+In most cases, QueryAttack still achieves high ASR, demonstrating its effectiveness on different datasets.
+C++ Python SQL Ensemble
+Gemini-1.5-flash 4.76 / 92.73% 4.74 / 89.09% 4.33 / 73.64% 4.84 / 94.55%
+GPT-4-1106 4.60 / 80.00% 4.38 / 72.73% 4.32 / 76.36% 4.65 / 83.64%
+DeepSeek-R1 4.81 / 88.18% 4.72 / 88.18% 4.71 / 87.27% 4.92 / 93.64%
+Llama-3.1-70B 4.38 / 76.37% 3.98 / 64.55% 2.09 / 14.55% 4.51 / 82.73%
+Table 2: QueryAttack’s attack performance on HEx-PHI (Zou et al., 2023a).
 QueryAttack remains effective when facing reasoning-enhanced models.
@@
 These results indicate that CoT Reasoning-enhanced models may have the potential to defend against QueryAttack, but QueryAttack still maintains a considerable ASR.
+Figure 4: Performance of different language styles.
 4.3 Ablation and Analysis
@@
 This may be because these two template styles closely resemble the structure of natural language, making them more likely to trigger existing defenses.
-Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reachs 76.3%, while the Ensemble configuration achieves 92.9%.
-Figure 4: Performance of different language styles.
-Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
-ASR HS ASR HS ASR HS ASR HS
-Paraphrase 94% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ 4.16 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ 68% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ 90% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.86 ${\color[rgb]{1,0,0}\uparrow 0.14}$
-Rand-insert 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 86% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.76 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ 66% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.46 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
-Rand-swap 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.88 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ 54% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ 4.12 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ 70% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
-Rand-patch 100% ${\color[rgb]{1,0,0}-0\%}$ 5.00 ${\color[rgb]{1,0,0}-0.00}$ 94% ${\color[rgb]{1,0,0}\uparrow 2\%}$ 4.90 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 64% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ 4.48 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ 72% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ 4.50 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
-Ours 36% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ 3.56 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ 28% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ 3.10 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ 76% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ 4.68 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ 34% ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ 3.38 ${\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
-Table 2: QueryAttack’s average ASR / HS against defense baselines. The differences between the defense and no defense are indicated by arrows.
+Despite this, under the Top 1 configuration, the ASR on Llama-3.1-70B still reaches 76.3%, while the Ensemble configuration achieves 92.9%.
+3.1-8B 3.1-70B 3.2-1B 3.2-3B 3.2-11B
+ASR 88.89% 92.91% 94.83% 81.61% 88.89%
+RR 2.87% 4.41% 3.45% 2.11% 3.26%
+Table 3: QueryAttack’s attack success rate and refusal rate on Llama 3.1 and 3.2 series models with different parameter sizes, under the Ensemble configuration.
 Larger models do not provide better defense against QueryAttack.
@@
 Therefore, exploring more robust defense methods is necessary.
-3.1-8B 3.1-70B 3.2-1B 3.2-3B 3.2-11B
-ASR 88.89% 92.91% 94.83% 81.61% 88.89%
-RR 2.87% 4.41% 3.45% 2.11% 3.26%
-Table 3: QueryAttack’s attack success rate and refusal rate on Llama 3.1 and 3.2 series models with different parameter sizes, under the Ensemble configuration.
+Figure 5: The attention score of natural language malicious query and QueryAttack.
 Attention score distribution of natural language malicious queries and QueryAttack.
@@
 In contrast, when applying query attack, LLMs mainly allocate attention to the term “method” and “WHERE NAME = ”, leading to a higher possibility to jailbreak.
-Figure 5: The attention score of natural language malicious query and QueryAttack.
+Gemini-flash GPT-4-1106 GPT-3.5 Llama-3.1-8B
+ASR HS ASR HS ASR HS ASR HS
+Paraphrase $94\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.88_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.12}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 20\%}$ $4.16_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.76}$ $68\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.56_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.14}$ $90\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.86_{\color[rgb]{1,0,0}\uparrow 0.14}$
+Rand-insert $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $86\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.76_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.16}$ $66\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ $4.48_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.46_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.26}$
+Rand-swap $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $94\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.88_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.04}$ $54\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 28\%}$ $4.12_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.58}$ $70\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 16\%}$ $4.50_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
+Rand-patch $100\%_{\color[rgb]{1,0,0}-0\%}$ $5.00_{\color[rgb]{1,0,0}-0.00}$ $94\%_{\color[rgb]{1,0,0}\uparrow 2\%}$ $4.90_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ $64\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 18\%}$ $4.48_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$ $72\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 14\%}$ $4.50_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.22}$
+Ours $36\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 74\%}$ $3.56_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.44}$ $28\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 64\%}$ $3.10_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.82}$ $76\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 6\%}$ $4.68_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 0.02}$ $34\%_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 52\%}$ $3.38_{\color[rgb]{0.1328125,0.546875,0.1328125}\downarrow 1.34}$
+Table 4: QueryAttack’s average ASR / HS against defense baselines.
+The differences between the defense and no defense are indicated by arrows.
+Shaded ellipses represents the 95% confidence interval (CI) for each data distribution.
+Embedding Differences Between Natural Language and QueryAttack.
+In addition to the analysis of attention mechanism, we further discuss the embedding differences between structured non-natural query languages and natural language in LLMs.
+Specifically, we use the embedding layer of Llama-3.2-1B (Meta, 2024) and OpenAI’s text-embedding-3-large model (Neelakantan et al., 2022) to embed both structured non-natural queries and their corresponding natural language queries.
+We randomly sampled 50 samples from AdvBench (Zou et al., 2023b) and compared the embedding differences between the original natural language queries and structured non-natural queries based on Java and C, and then analyze and visualize their differences using t-SNE, with the results presented in Figure 6.
+The results demonstrate a significant divergence in embeddings between structured non-natural query languages and natural language in LLMs, suggesting that QueryAttack could bypass deployed safety alignment mechanisms with a high probability.
+(a) Llama-3.2-1B (C and Java)
+(b) OpenAI-Emb-Lg (C and Java)
+Figure 6: The t-SNE visualizations of the embedding representations of original queries and QueryAttack using Llama-3.2-1B and text-embedding-3-large.
 4.4 Discussion about Countermeasures
-We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix B.2.
+We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix C.2.
 Paraphrase (Jain et al., 2023a): A defense method that reduces the ASRs by reconstructing inputs while preserving natural semantics.
@@
 The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense.
-The detailed prompt can be found in the Appendix B.2.
+The detailed prompt can be found in the Appendix C.2.
 We test the effectiveness of these countermeasures against QueryAttack on four models: Gemini-flash, GPT-3.5, Llama-3.1-8B, and GPT-4-1106.
@@
 Although QueryAttack does not encrypt the outputs, our extensive evaluation shows that it still effectively bypasses the defenses of mainstream LLMs and can withstand common defense methods.
-Besides, to defend against QueryAttack, we develop a tailored defense strategy.
-Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
+Besides, to defend against QueryAttack, we develop a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
+Acknowledgments
+We thank the anonymous reviewers for their feedback on this work.
+This work is supported by the Major Key Project of PCL under grant No.
+PCL2023A06, Shenzhen R&D Program under grant NO.
+KJZD20230923114059020, and in part by Shenzhen Key Laboratory of Software Defined Networking under Grant ZDSYS20140509172959989.
 Limitations
@@
 We aim to highlight the potential vulnerabilities about QueryAttack and encourage practitioners to prioritize the development of robust security measures and trustworthy deployment practices.
-Appendix A Attack Examples of QueryAttack
-A.1 Success Examples
-Figure 6 illustrates a simplified attack process of QueryAttack. The adversary first utilizes query learning in context. Then, the translated query is used to launch the attack.
-Figure 6: An attack example of QueryAttack using a C++ style template.
-A.2 Prompts of Query Learning
+Appendix A Evaluation about the Capability of LLMs in Understanding and Generating Encrypted Content
+To evaluate the capability of LLMs in understanding and generating encrypted content, we adopt the prompt used in CipherChat (Yuan et al., 2024) for Caesar ciphers and construct a benign request (as malicious ones might trigger defense mechanisms): “Please tell me a story for children.
+This story should include a big bad wolf and a little girl.” We then translate it into a Caesar cipher with a shift of 3 and submitted it to mainstream LLMs.
+We evaluate the models’ responses from two perspectives: (1) whether the model could accurately understand the encrypted text, and (2) whether it could generate a high-quality response that meets user expectations (i.e., free of readability issues and logically coherent).
+The results are shown in Table 5.
+Model GPT-4o Gemini-1.5-Pro Llama-3.1-8B DeepSeek-R1 GPT-4.5
+Understanding
+Correctness
+Yes No No Yes Yes
+Generation
+Quality
+No No No No Yes
+Table 5: Capability of LLMs in Understanding and Generating Encrypted Text.
+As an example, here is a response from GPT-4o (after translated into natural language): “It looks like a fair| tale!
+Here’s a nice mi[ of a big bad wolf and a little girl: 1.
+Big bad wolf = This could be a reference to a truck, an anquatic and dangerods one.
+2.
+Little girl = Ma|be a little girl is a compact temple and a late gravitation surge.
+You might be pointing to a scene from a sci-fi movie or book—specificall| a polot involving a large and dangerods poawer fuell incorporated into a temple with a compact surge s
+stem. Would you like me to e[plore a narrative based on this ciffer”
+Appendix B Attack Examples of QueryAttack
+B.1 Prompts of Query Learning
 Figure 7 and Figure 8 present the prompts used for few-shot and zero-shot learning in SQL style.
@@
 Figure 8: The few-shot prompts for Query Understanding (C++ style).
-Appendix B Experimental Details
-B.1 Human Evaluation on GPT-4 Judge
+B.2 Success Examples
+Figure 9 illustrates a simplified attack process of QueryAttack. The adversary first utilizes query learning in context. Then, the translated query is used to launch the attack.
+Figure 9: An attack example of QueryAttack using a C++ style template.
+Appendix C Experimental Details
+C.1 Human Evaluation on GPT-4 Judge
 To confirm the effectiveness of GPT evaluation, we conduct an experiment with human evaluators to assess the responses of LLMs.
@@
 The result demonstrates the effectiveness of GPT-4 based evaluation.
-B.2 Defense Settings
+C.2 Defense Settings
 Paraphrase.
@@
 Figure 10: Prompt for cross-lingual alignment based defense.
-B.3 Ethic-related Disclaimer Removal
+C.3 Ethic-related Disclaimer Removal
 Figure 11 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 7.
@@
 The response is clearly malicious, containing a lot of harmful knowledge, but includes concise mitigation measures as required and provides educational disclaimers at the beginning and end (highlighted in purple).
-Appendix C Prompts for Extracting Query Components
+Appendix D Prompts for Extracting Query Components
 We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process.